OpenStudy (anonymous):

Kristos is using a 14 foot ladder to reach an attic vent on the side of his house. He places the ladder 4 feet from the base of teh house and leans it against the house until it just touches the vent. What is the height of the attic vent? a. 10 ft b. 13.4 feet c. 14.6 feet d. 18 feet Please help...thanks.

OpenStudy (anonymous):

use trig for this

OpenStudy (anonymous):

i don't understand

OpenStudy (anonymous):

If the angle that the ladder makes with the ground is theta then cosine of this angle is equal to he ratio 4/14

OpenStudy (anonymous):

then take the inverse cosine to find the angle made with the ladder and the ground

OpenStudy (anonymous):



OpenStudy (anonymous):

use Pythagorus theorem. c^2=a^2+b^2 14^2-4^2=b^2 196-16=180 b=sqrt(180) 13.4

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

you're welcome.